# SciAssess: Benchmarking LLM Proficiency in Scientific Literature Analysis

### v3

@@ 3.2.2 Performance on Multimodal Contents
Table 4: Performances on multimodal contents. Orange and green indicate the best in closed and open source LLMs, respectively.
Modality Question Type Metric GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command $\text{R}^{\text{+}}$
Text Only Multiple Choice Accuracy 0.77 0.74 0.59 0.76 0.73 0.69 0.76 0.65 0.74 0.77 0.55
Text Extraction Recall 0.85 0.86 0.81 0.81 0.85 0.79 0.82 0.81 0.82 0.78 0.76
True/False Accuracy 0.74 0.72 0.46 0.76 0.74 0.74 0.70 0.54 0.69 0.66 0.59
Average Rank 3.14 4.86 8.29 4.32 5.32 7.18 6.00 6.61 5.89 4.71 9.68
Chart Multiple Choice Accuracy 0.56 0.54 0.30 0.52 0.39 0.59 0.65 0.56 0.50 0.56 0.52
Average Rank 4.88 6.62 10.50 6.00 8.62 4.62 2.38 4.75 6.62 5.12 5.88
Chemical Reaction Multiple Choice Accuracy 0.62 0.66 0.45 0.40 0.56 0.54 0.52 0.56 0.46 0.52 0.45
Average Rank 4.00 2.25 8.25 9.00 6.75 4.25 5.50 6.50 7.50 5.25 6.75
Molecular Structure Molecule Generation Mol Similarity 0.34 0.34 0.22 0.26 0.36 0.17 0.35 0.26 0.24 0.19 0.20
Table Extraction Table Accuracy 0.20 0.22 0.11 0.04 0.29 0.15 0.07 0.16 0.15 0.23 0.14
True/False Accuracy 0.58 0.70 0.50 0.46 0.48 0.56 0.64 0.68 0.46 0.46 0.46
Average Rank 3.40 3.80 8.70 8.40 2.80 6.00 5.30 4.60 8.30 6.80 7.90
Table Multiple Choice Accuracy 0.89 0.79 0.37 0.67 0.87 0.71 0.88 0.46 0.72 0.62 0.45
Table Extraction Table Accuracy 0.40 0.41 0.22 0.25 0.42 0.30 0.28 0.29 0.36 0.39 0.19
Average Rank 2.33 3.33 9.33 7.92 2.67 7.00 7.08 7.00 5.50 4.67 9.17
Table 4 summarizes the performance of LLMs on multimodal content tasks, encompassing text-only, chart, molecular structure, chemical reaction, and table modalities. For each modality, performances are averaged over different question types. We observe the following: (1) Text-only tasks: GPT-4o achieves the highest accuracy for multiple-choice questions (0.77), leading to the best average rank (3.14). (2) Chart tasks: Gemini exhibit the highest accuracy (0.65), resulting in average ranks of 2.38. (3) Chemical reaction tasks: GPT-4 stands out with the best average rank (2.25), indicating high accuracy in multiple-choice questions (0.66). (3) Molecule tasks: Claude3 excels with average ranks of 2.80, particularly in molecule generation and table extraction tasks. (5) Table tasks: GPT-4o lead with the highest multiple-choice accuracy (0.89) and average rank of 2.33.
Overall, GPT-4o consistently ranks as a top performer across all modalities, particularly excelling in text-only and table tasks. GPT-4 and Claude3 also demonstrate strong performance, especially in molecular structure and chemical reaction tasks. Doubao and Gemini are competitive in specific modalities, such as chart tasks and molecule-related evaluations.
Based on these observations, we suggest the following recommendations: (1) For text-only tasks, GPT-4o and GPT-4 are highly recommended due to their superior accuracy and ranking. (2) For molecular structure and chemical reaction tasks, Claude3 and GPT-4 excel, making them suitable for such specialized applications. (3) For table tasks, Claude and GPT-4o are the preferred models, given their high accuracy in multiple-choice and table extraction tasks.
3.2.3 Performance w/ and w/o Contexts
Table 5: Performance with and without contexts. Orange and green indicate the best in closed and open source LLMs, respectively.
@@ 3.2.3 Performance w/ and w/o Contexts
Modality Question Type Metric GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command R+
W/o Context Molecule Generation Mol Similarity 0.64 0.65 0.43 0.50 0.68 0.24 0.53 0.49 0.47 0.38 0.38
Multiple Choice Accuracy 0.77 0.70 0.59 0.73 0.69 0.73 0.73 0.71 0.72 0.77 0.60
Text Extraction Recall 0.87 0.87 0.86 0.85 0.86 0.82 0.83 0.85 0.84 0.85 0.81
Average Rank 2.95 5.35 7.70 4.65 5.70 8.00 6.75 5.25 6.10 4.10 9.45
With Context Molecule Generation Mol Similarity 0.04 0.02 0.00 0.02 0.04 0.09 0.17 0.03 0.01 0.00 0.03
Multiple Choice Accuracy 0.67 0.66 0.42 0.60 0.59 0.61 0.70 0.55 0.59 0.62 0.49
Table Extraction Table Accuracy 0.40 0.41 0.22 0.25 0.42 0.30 0.28 0.29 0.36 0.39 0.19
Text Extraction Recall 0.72 0.79 0.53 0.61 0.79 0.62 0.77 0.60 0.69 0.41 0.48
True/False Accuracy 0.66 0.71 0.48 0.61 0.61 0.65 0.67 0.61 0.57 0.56 0.52
Average Rank 3.47 3.97 9.37 6.63 4.92 5.66 4.61 6.68 6.39 5.95 8.34
Table 5 presents the performance of LLMs on different tasks with and without context. We observe the following: (1) GPT-4o consistently ranks highly both with and without context, demonstrating robustness across different modalities. (2) Gemini shows significant improvement with context, particularly in molecule generation and text extraction. This could be due to their proprietary PDF parsing technology being more advanced than the open-source PyPDF. (3) GPT-4 and Claude3 maintain strong performance, especially in context-dependent tasks. (4) GPT-3.5 and Command $\text{R}^{\text{+}}$ generally underperform across both conditions, indicating a need for optimization and improvement for science literature analysis.
Based on these observations, we suggest the following: (1) For tasks without contexts, GPT-4o is recommended due to its superior performance and consistency. (2) For tasks with contexts, Gemini and GPT-4o are excellent choices.
@@ 4 Related Work
General benchmarks for LLMs. Large language models (LLMs) are evaluated across a variety of benchmarks to comprehensively assess their capabilities. Some benchmarks, such as MMLU [12], MMLU-pro [22], CMMLU [13], and Xiezhi [14], are instrumental in evaluating models’ world knowledge across diverse domains. For reasoning capabilities, benchmarks like GSM8k [23] and BBH [24] provide rigorous assessments of models’ problem-solving and logical reasoning skills. In the realm of programming, benchmarks such as HumanEval [25] and MBPP [26] serve as popular testbeds for evaluating models’ coding proficiency. Open-ended generation tasks, essential for measuring creative and coherent text production, are evaluated using benchmarks such as AlpacaEval [27]. Additionally, TruthfulQA [28] and HaluEval [29] are pivotal in assessing the veracity of models’ outputs, ensuring their alignment with factual information.
Although some general benchmarks include a subset of science subjects, they mostly focus on Memorization (L1) and often overlook higher-level abilities such as Comprehension (L2) and Analysis & Reasoning (L3). Furthermore, these benchmarks lack context-involved scientific literature tasks, for example, understanding and reasoning over a scientific paper.
Scientific literature benchmarks. Prior works have made significant strides in developing LLM benchmarks aimed at enhancing and assessing the understanding of scientific literature. In the biomedical domain, notable efforts include BLUE [30], which provides a set of tasks for evaluating models on various aspects of biomedical text-mining. Building on this, BLURB [31] offers an extensive collection of datasets to further refine model performance in this specialized field. More recently, InBoXBART [32] has been introduced, focusing on integrating information across multiple biomedical documents.
Nevertheless, when it comes to scientific literature analysis, these benchmarks are confined to the biomedical domain, neglecting other fields such as chemistry and materials science. Furthermore, they lack adequate tasks for interpreting multi-modal content (e.g., molecular structures and tables), which are common in real-world scientific literature.
@@ 5 Conclusion and Future Work
SciAssess rigorously assesses the capabilities of LLMs in the domain of scientific literature analysis. It is built upon fundamental science and extends to four specialized areas: Alloy Materials, Biomedicine, Drug Discovery, and Organic Materials. The benchmark focuses on assessing LLMs’ core competencies in Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3) within the context of scientific literature analysis. Through detailed evaluations of 11 LLMs, SciAssess highlights their strengths and identifies areas needing improvement across various ability levels, content modalities, and contextual scenarios. Additionally, we emphasize the urgent need for PDF parsing algorithms tailored to handle content of various modalities, such as molecular structures and chemical reactions. We hope that SciAssess supports the ongoing development of LLMs in scientific literature analysis. Looking ahead, we plan to broaden the range of scientific domains included in SciAssess and incorporate more vertical domains. These enhancements aim to improve the benchmark’s utility and efficacy, providing clearer guidance and fostering the advancement of LLMs in scientific literature analysis.
@@
While SciAssess provides a comprehensive and valuable benchmarking suite across five primary domains—fundamental science, alloy materials, biomedicine, drug discovery, and organic materials—there are several limitations to consider. Firstly, the scope of SciAssess is currently constrained to these five domains, with potential future extensions to other vertical domains such as physics and engineering.
Secondly, the creation and curation of high-quality, domain-specific training data are essential for the effective evaluation and improvement of LLMs. However, due to the high costs associated with manual labeling, SciAssess does not provide additional training data for these tasks. This absence of supplementary data can limit the ability of researchers to fine-tune and enhance LLMs specifically for the tasks included in SciAssess. Consequently, the benchmark results might reflect the inherent capabilities of the models rather than their optimized performance for each specific domain.
@@
Appendix
@@ Appendix A Question Type
Five types of questions, as illustrated in Figure 5 are devised to evaluate the models: True/False questions, Multiple Choice questions, Open-ended Generation, Constrained Generation, and Table Extraction. Each question type is accompanied by a detailed description and representative examples, along with the corresponding metrics used for assessment. For convenience, the input in each example is simplified, and its instruction is omitted.
@@
Appendix B Prompt Template
We design following prompt template for scientific literature analysis assistant:
Prompt Template Role setting and task description: You are a highly intelligent assistant who answers the following multiple choice question correctly. Only write the answer down. ICL: Question: xxx Answer: xxx Question: xxx Answer: xxx Question: Predict the number of lines in the EPR spectrum of a solution of 13C-labelled methyl radical (13CH3•), assuming the lines do not overlap. a) 4 b) 3 c) 6 d) 24
@@ Appendix C Task Prompt
C.1 Fundamental Science
Prompt System Message: You are a highly intelligent assistant who answers the following multiple choice question correctly. Only write the answer down. User Message: Predict the number of lines in the EPR spectrum of a solution of 13C-labelled methyl radical (13CH3•), assuming the lines do not overlap. a) 4 b) 3 c) 6 d) 24 Expected Answer: a) 4
C.2 Alloy Chart QA
Prompt System Message: You are an expert in the field of Alloy Materials. You are a highly intelligent alloy researcher who answers the following multiple choice question correctly. Only write the closest option from the listed options, such as ’b) 2045.1’. Don’t change the option values. User Message: In FIG. 9, what is the closest maximum pore area (with unit $\text{um}^{2}$) of batches $\text{A}_{Ti}$? a) 54000 b) 45000 c) 5000 d) 10000 Expected Answer: a) 54000
C.3 Composition Extraction
An alloy composition table example is shown as following:
Table 6: Alloy composition example.
AlloyName Composition Composition Composition Composition Composition Composition
/ C Cr Cu Fe Mn Mo
LeanDSS 0.014 % 20.85 % 0.09 % 73.38 % 1.49 % 0.30 %
StandardDSS 0.012 % 22.46 % 0.17 % 69.94 % 1.81 % 3.07 %
SuperDSS 0.013 % 24.98 % 0.20 % 63.41 % 0.48 % 4.03 %
Prompt User Message: You are an expert in the field of Alloy Materials. Please give a complete list of alloy names and compositions of all alloys in this paper. If there is no alloy composition element ratio in the text, try to extract the element ratio from the alloy name from the perspective of alloy experts. Output in csv format with multiindex (2 headers), The names in first header are ’AlloyName’ and ’Composition’ forcely. The names in second header are element names of alloy. Starting on the third row, list the alloy names and their corresponding element content. Based on the number of reference commas, the element name corresponds to the content. Please write units not in header but in value like "50 wt.%","30 at.%". Output the data strictly in the CSV format shown below and exclude any other content. Example format: “‘csv AlloyName,Composition,Composition,Composition nan,Fe,Co,Al Fe70Co15Al3,70 wt.%,15 wt.%,3 wt.Fe70Co18,70 wt.%,18 wt.%,nan ”’
C.4 Temperature Extraction
Examples of process paragraphs [33]:
We prompt the model with the following:
Prompt System Message: You are an expert in the field of Alloy Materials. You are a highly intelligent alloy researcher who answers the following multiple choice question correctly. Only write the option down, such as ’a) 925 C’, without explanation. User Message: In the upper paper, what is the maximum temperature of the heat treatment process for all alloys? a) 925 C b) 650 C c) 700 C d) 800 C Expected Answer: a) 925 C
C.5 Sample Differentiation
The following example is process paragraphs where the sample are treated by different processes [34]:
We prompt the model with the following:
Prompt System Message: You are an expert in the field of Alloy Materials. Please answer the following multiple choice question correctly. Only write the option (e.g., ’a) 3’) without explanation. User Message: Materials with the same components but processed through different techniques are considered as different alloys because they possess distinct properties. In the upper paper, please provide a count of all the alloys proposed and discussed by the authors? a) 2 b) 0 c) 3 d) 1 Expected Answer: d) 1
C.6 Treatment Sequence
Prompt System Message: You are an expert in the field of Alloy Materials. You are a specialist in the domain of heat treatment processes, such as homogenization, annealing, aging, solution treatment, quenching, and tempering, among others. Answer the following question with "Yes" or "No". User Message: In the upper paper, is the processing heat treatment technique before the thermal treatment at 925 C called arc melting? Expected Answer: Yes
C.7 Chemical Entities Recognition
@@
Prompt System Message: You are an expert in the field of Biomedical. I’ll give you the abstract of literature. Please identify all the compound entities in the abstract, and just give me a list of the compound entities you recognized without any other interpretation: (compound 1), (compound 2), (compound 3) User Message: In unanesthetized, spontaneously hypertensive rats the decrease in blood pressure and heart rate produced by intravenous clonidine, 5 to 20 micrograms/kg, was inhibited or reversed by nalozone, 0.2 to 2 mg/kg. The hypotensive effect of 100 mg/kg alpha-methyldopa was also partially reversed by naloxone. Naloxone alone did not affect either blood pressure or heart rate. In brain membranes from spontaneously hypertensive rats clonidine, 10(-8) to 10(-5) M, did not influence stereoselective binding of [3H]-naloxone (8 nM), and naloxone, 10(-8) to 10(-4) M, did not influence clonidine-suppressible binding of [3H]-dihydroergocryptine (1 nM). These findings indicate that in spontaneously hypertensive rats the effects of central alpha-adrenoceptor stimulation involve activation of opiate receptors. As naloxone and clonidine do not appear to interact with the same receptor site, the observed functional antagonism suggests the release of an endogenous opiate by clonidine or alpha-methyldopa and the possible role of the opiate in the central control of sympathetic tone. Expected Answer: (naloxone), (clonidine), (nalozone), (alpha-methyldopa)
C.8 Disease Entities Recognition
Prompt System Message: You are an expert in the field of Biomedical. You are a biologist AI. I’ll give you the abstract of literature. Please identify all the disease entities in the abstract, and please only give me a list of the disease entities you recognized without any other interpretation: (disease 1), (disease 2), (disease 3) User Message: In unanesthetized, spontaneously hypertensive rats the decrease in blood pressure and heart rate produced by intravenous clonidine, 5 to 20 micrograms/kg, was inhibited or reversed by nalozone, 0.2 to 2 mg/kg. The hypotensive effect of 100 mg/kg alpha-methyldopa was also partially reversed by naloxone. Naloxone alone did not affect either blood pressure or heart rate. In brain membranes from spontaneously hypertensive rats clonidine, 10(-8) to 10(-5) M, did not influence stereoselective binding of [3H]-naloxone (8 nM), and naloxone, 10(-8) to 10(-4) M, did not influence clonidine-suppressible binding of [3H]-dihydroergocryptine (1 nM). These findings indicate that in spontaneously hypertensive rats the effects of central alpha-adrenoceptor stimulation involve activation of opiate receptors. As naloxone and clonidine do not appear to interact with the same receptor site, the observed functional antagonism suggests the release of an endogenous opiate by clonidine or alpha-methyldopa and the possible role of the opiate in the central control of sympathetic tone. Expected Answer: (hypertensive), (hypotensive)
C.9 Compound Disease Recognition
Examples of process text:

### v5

@@ 3.2.2 Performance on Multimodal Contents
Modality Question Type Metric o1 GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3.1 Qwen2.5 Mixtral
Text Only Multiple Choice Accuracy 0.798 0.788 0.740 0.423 0.707 0.523 0.594 0.737 0.724 0.729 0.540
Text Extraction F1-score 0.781 0.675 0.707 0.591 0.690 0.686 0.691 0.705 0.718 0.677 0.655
True/False Accuracy 0.624 0.594 0.678 0.485 0.683 0.480 0.658 0.634 0.614 0.658 0.673
Average Rank 3.000 6.667 2.667 10.667 4.667 9.333 5.833 4.667 5.333 5.833 7.333
Chart Multiple Choice Accuracy 0.696 0.590 0.427 0.428 0.496 0.451 0.464 0.604 0.536 0.538 0.522
Average Rank 1.000 3.000 11.000 10.000 7.000 9.000 8.000 2.000 5.000 4.000 6.000
Reaction Multiple Choice Accuracy 0.620 0.539 0.357 0.250 0.458 0.188 0.340 0.513 0.420 0.440 0.252
Average Rank 1.000 2.000 7.000 10.000 4.000 11.000 8.000 3.000 6.000 5.000 9.000
Mol. Table Extraction Recall 0.231 0.270 0.266 0.168 0.112 0.050 0.247 0.236 0.305 0.285 0.202
Mol. Generation Mol. Similarity 0.394 0.407 0.388 0.273 0.262 0.282 0.335 0.447 0.284 0.290 0.298
True/False Accuracy 0.840 0.600 0.620 0.440 0.480 0.640 0.560 0.580 0.520 0.600 0.460
Average Rank 3.667 3.167 3.667 10.000 10.000 7.333 5.667 4.333 5.667 4.500 8.000
Table Multiple Choice Accuracy 0.925 0.855 0.810 0.305 0.765 0.435 0.745 0.765 0.755 0.785 0.455
Table Extraction Recall 0.397 0.441 0.437 0.299 0.365 0.307 0.362 0.376 0.441 0.426 0.294
Average Rank 3.000 1.750 3.000 10.500 6.250 9.500 8.000 5.750 4.250 4.000 10.000
Table 4: Performance on multimodal contents.
Table 4 summarizes the performance of LLMs on multimodal content tasks. For each modality, performances are averaged over different question types. We observe the following: (1) Text-only tasks: GPT-4 achieves the highest average rank (2.00). (2) Chart tasks: OpenAI-o1 exhibit the highest accuracy (0.696). (3) Chemical reaction tasks: OpenAI-o1 stands out with high accuracy in multiple-choice questions (0.62). (3) Molecule tasks: GPT-4o excels with average ranks of 3.17, particularly in table extraction task. (5) Table tasks: GPT-4o lead with the highest table extraction recall (0.44).
Overall, OpenAI models consistently rank as top performers across most modalities. Gemini also demonstrate strong performance, especially in molecule generation tasks.
Based on these observations, we suggest the following recommendations: (1) For text-only tasks, OpenAI-o1 and GPT-4o are highly recommended due to their superior accuracy and ranking. (2) For chart and chemical reaction tasks, OpenAI-o1 excels, making it suitable for such specialized applications. (3) For molecule structure and tabular tasks, GPT-4o is the preferred model, given it remarkable performance.
@@ 4 Related Work
General benchmarks for LLMs. LLMs are evaluated across a variety of benchmarks to comprehensively assess their capabilities. Some benchmarks, such as MMLU Hendrycks et al. (2021), MMLU-pro Wang et al. (2024), CMMLU Li et al. (2023a), and Xiezhi Gu et al. (2024), are instrumental in evaluating models’ world knowledge across diverse domains. For reasoning capabilities, benchmarks like GSM8k Cobbe et al. (2021) and BBH Suzgun et al. (2023b) provide rigorous assessments of models’ problem-solving and logical reasoning skills. In the realm of programming, benchmarks such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) serve as popular testbeds for evaluating models’ coding proficiency. Additionally, TruthfulQA Lin et al. (2022) and HaluEval Li et al. (2023b) are pivotal in assessing the veracity of models’ outputs, ensuring their alignment with factual information.
Although some general benchmarks include a subset of science subjects, they mostly focus on Memorization (L1) and often overlook higher-level abilities such as Comprehension (L2) and Analysis & Reasoning (L3). Furthermore, these benchmarks lack context-involved tasks, for example, understanding and reasoning over a scientific paper.
Scientific literature benchmarks. Prior works have made significant strides in developing LLM benchmarks to assess the understanding of scientific literature. In the biomedical domain, notable efforts include BLUE Peng et al. (2019), which provides a set of tasks for evaluating models on various aspects of biomedical text-mining. Building on this, BLURB Gu et al. (2021) offers an extensive collection of datasets to further refine model performance in this specialized field. More recently, InBoXBART Parmar et al. (2022) has been introduced, focusing on integrating information across multiple biomedical documents. SciRIFF Wadden et al. (2024) is designed to extract and synthesize information from research literature across various scientific disciplines.
Compared with existing scientific literature benchmarks, SciAssess focuses more on tasks for interpreting multi-modal content (e.g., molecular structures and tables), which are common in scientific literature. Moreover, it features a real-world application scenarios that LLMs digest parsed PDF contents with parsing errors.
@@ 5 Conclusion and Future Work
SciAssess rigorously assesses the capabilities of LLMs for scientific literature analysis. It focuses four specialized areas: biology, chemistry, material, and medicine. The benchmark focuses on assessing LLMs’ core competencies in Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3) within the context of scientific literature analysis. Through detailed evaluations of 11 LLMs, SciAssess highlights their strengths and identifies areas needing improvement across various ability levels, content modalities, and contextual scenarios. Additionally, we emphasize the urgent need for PDF parsing algorithms tailored to handle content of various modalities, such as molecular structures and chemical reactions. We hope that SciAssess supports the ongoing development of LLMs in scientific literature analysis. Looking ahead, we plan to broaden the range of scientific domains included in SciAssess and incorporate more vertical domains. These enhancements aim to improve the benchmark’s utility and efficacy, providing clearer guidance and fostering the advancement of LLMs in scientific literature analysis.
@@
While SciAssess provides a comprehensive and valuable benchmarking suite across four primary domains – biology, chemistry, material, and medicine – there are several limitations to consider. Firstly, the scope of SciAssess is currently constrained to these four domains, with potential future extensions to other vertical domains such as physics and engineering.
Secondly, the creation and curation of high-quality, domain-specific training data are essential for the effective evaluation and improvement of LLMs. However, due to the high cost associated with manual labeling, SciAssess does not provide additional training data for these tasks. This absence of supplementary data can limit the ability of researchers to fine-tune and enhance LLMs specifically for the tasks included in SciAssess. Consequently, the benchmark results might reflect the inherent capabilities of the models rather than their optimized performance for each specific domain.
@@
Broader Impact
Our work on benchmarking scientific literature analysis aligns with the scope of existing LLM benchmarks such as MMLU-pro. This paper represents progress in calibrating LLMs for specific domains, thereby amplifying the impacts that LLM benchmarks have had (and will continue to have) on the broader world. Additionally, we have not identified any ethical concerns or potential risks associated with this work.
@@ Appendix A Question Type
Five types of questions, as illustrated in Figure 5 are devised to evaluate the models. Each question type is accompanied by a detailed description and representative examples, along with the corresponding metrics used for assessment. For convenience, the input in each example is simplified, and its instruction is omitted.
@@
Appendix B General Prompt Template
We design following general prompt template for scientific literature analysis. It consists of: a system message defining the role of the assistant, the task description, some optional few-shot examples, and a user prompt of the question.
Prompt Template Role setting and task description: You are a highly intelligent assistant who answers the following multiple choice question correctly. Few-shot examples: Question: <question 1> Answer: <answer 1> Question: <question n> Answer: <answer n> Question: Predict the number of lines in the EPR spectrum of a solution of 13C-labelled methyl radical (13CH3•), assuming the lines do not overlap. a) 4 b) 3 c) 6 d) 24
@@ Appendix C Task Prompt
In this section, we detail the prompt templates for all tasks in SciAssess benchmark. We will introduce these templates in the following order: Biology (Section C.1), Chemistry (Section C.2), Material (Section C.3) and Medicine (Section C.4).
C.1 Biology
C.1.1 MMLU-Pro-Biology
Prompt System Message: You are a highly intelligent assistant who answers the following multiple choice question correctly. Use chain of thought reasoning and provide reasoning before selecting the correct answer (e.g., a)xxx, or b)xxx). Format: Reasoning: [Reasoning] Answer: [Answer] User Message: Which of the following would most likely provide examples of mitotic cell divisions? a) cross section of muscle tissue b) longitudinal section of a shoot tip c) longitudinal section of a leaf vein d) cross section of a fruit e) cross section of a leaf f) longitudinal section of a petal g) longitudinal section of a seed h) cross section of an anther (site of pollen production in a flower) Expected Answer: b) longitudinal section of a shoot tip
C.1.2 Biology Chart QA
The analysis and understanding of biological properties, compositions, and processing techniques are critical for the discovery and development in life sciences. Often, this information is presented in charts, making it essential to extract and integrate such information with textual data. To assess the retrieval capabilities of models in the context of biological chart information, we have designed multiple-choice questions.
Prompt System Message: You are an expert in the field of Biomedical. You are a highly intelligent biology scientist who answers the following multiple choice question correctly. Use chain of thought reasoning and provide reasoning before selecting the correct answer (e.g., a)xxx, or b)xxx). Format: Reasoning: [Reasoning] Answer: [Answer] User Message: In Figure 3, which has a higher accurate score, with the graph encoder or without? a) with graph encoder b) w/o graph encoder Expected Answer: a) with graph encoder
C.1.3 Chemical Entities Recognition
This task involves recognizing chemical entity names using data from B5CDR Wei et al. (2016) and additional expert-annotated data. It evaluates the performance of LLMs in identifying complex drug names. The prompt template is as follows.
@@ C.1.3 Chemical Entities Recognition
Prompt System Message: You are an expert in the field of Biomedical. I’ll give you the abstract of literature. Please use chain of thought reasoning to identify all the compound entities in the abstract. First, analyze the abstract step by step, explaining your reasoning for identifying each compound entity. Then, provide a final list of the compound entities you recognized in the format: (compound 1), (compound 2), (compound 3). Format: Reasoning: [Reasoning] Answer: [List of identified compounds] User Message: In unanesthetized, spontaneously hypertensive rats the decrease in blood pressure and heart rate produced by intravenous clonidine, 5 to 20 micrograms/kg, was inhibited or reversed by nalozone, 0.2 to 2 mg/kg. The hypotensive effect of 100 mg/kg alpha-methyldopa was also partially reversed by naloxone. Naloxone alone did not affect either blood pressure or heart rate. In brain membranes from spontaneously hypertensive rats clonidine, 10(-8) to 10(-5) M, did not influence stereoselective binding of [3H]-naloxone (8 nM), and naloxone, 10(-8) to 10(-4) M, did not influence clonidine-suppressible binding of [3H]-dihydroergocryptine (1 nM). These findings indicate that in spontaneously hypertensive rats the effects of central alpha-adrenoceptor stimulation involve activation of opiate receptors. <rest of the input>. Expected Answer: (naloxone), (clonidine), (nalozone), (alpha-methyldopa)
C.1.4 Compound Disease Recognition
Proposed in B5CDR Wei et al. (2016), this task evaluates the capability of LLMs to identify and understand associations between compounds and diseases. Examples of process text: